OpenStudy (anonymous):

wha is the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12).

OpenStudy (kropot72):

The slope is the distance between the y points divided by the distance between the x points = (12 - 7)/(3 - (-2)) = 5/5 = 1 The general equation is y = mx + b where m is the slope and b is the y-intercept. Substituting the values (3, 12) gives 12 = (1 * 3) + b 12 = 3 + b b = 12 - 3 = 9 The required equation is y = x + 9

OpenStudy (kropot72):

Do you understand?

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

You're welcome :)